QUESTION 9 Jimmy is the owner of the only gelato shop on North Stradbroke lsland. He sells the gelato in cones, with the marginal cost of each gelato cone being constant at $0.80 per cone. Below is a table of potential prices and corresponding quantities of gelato cones he could sell. Quantity (cones) 50 60 70 80 90 100 Price ($/cone) 6.80 6.20 5.60 5.00 4.40 3.80 What is Jimmys profit-maximising quantity of cones? Answer to the nearest whole number (with no decimal places)
Answer #1

Jimmy's Profit maximizing quantity of cones is where MR = MC

Given that MC = 0.8, so profit maximizing quantity of cones is where MR = MC = $0.80

Price Quantity Total revenue Marginal revenue
6.80 50 340 -----
6.20 60 372 (372 - 340) / ( 60 -50) = 3.2
5.60 70 392 (392 - 372) / ( 70 - 60) = 2.0
5.00 80 400 (400 - 392) / ( 80 - 70) = 0.8
4.40 90 396 -0.4
3.80 100 380 -1.6

According to the table, MR = MC = 0.8 where quantity of cones is 80,

So,

Jimmy's Profit maximizing quantity of cones is 80 cones.